OpenStudy (sally6sa):

A sweepstakes give away $1 000 000 prizes the first ticket drawn wins $5 , the second ticket drawn wins $15 , the third drawn wins $45 , and so on . 1) how many tickets can sweepstakes afford to draw ? 2) how much money is left after all the prizes are awarded ?

OpenStudy (shirleyarnett):

The sum of 5 * 3^k (from k=0 to n) <= 1 million. The sum of 3^k (from k=0 to n) <= 200,000 Floor[3^(n+1) / 2] <= 200,000 3^(anything) is odd, so Floor[3^(n+1) / 2] = [3^(n+1) - 1] / 2 3^(n+1) <= 400,001 Now solve for n.